OpenStudy (kyleandry):

Why isn't the (complex) dot product of 2 vectors (x and y) defined as sum_{k=0}^{n}(\Re(x_{k})*\Re(y_{k})+\Im(x_{k})*\Im(y_{k})) ?

OpenStudy (zzr0ck3r):

\(\sum_{k=0}^{n}(\Re(x_{k})*\Re(y_{k})+\Im(x_{k})*\Im(y_{k})) \) What does that mean?

OpenStudy (unklerhaukus):

\[%x_k = \Re(x_k)+\Im(x_k),\ \qquad y_k = \Re(y_k)+\Im(y_k)\\[2ex] x_k\cdot y_k=\Big(\Re(x_k)+\Im(x_k)\Big)\cdot\Big(\Re(y_k)+\Im(y_k)\Big)\\ \qquad\quad = \Re(x_k)\cdot\Re(y_k)+\Re(x_k)\cdot\Im(y_k)+\Im(x_k)\cdot\Re(y_k)+\Im(x_k)\cdot\Im(y_k)\]